# Beyond the Trade-off Curve: Multivariate and Advanced Risk-Utility Maps for Evaluating Anonymized and Synthetic Data

### v1

@@ Abstract
Keywords: risk-utility visualization, statistical disclosure control, data anonymization, synthetic data, principal component analysis, biplot, pareto optimality, multivariate visualization
@@ 1 Introduction
In microdata anonymization, modifications to the original dataset, such as suppression, generalization, perturbation, or synthetic data generation, are essential to reduce disclosure risk [20, 42]. However, these modifications inevitably result in a loss of information, potentially limiting the analytical value of the data. Effective anonymization therefore requires balancing the confidentiality gained from lowering disclosure risk with the preservation of data utility. A possibility for evaluating this trade-off visually is the risk–utility (R-U) confidentiality map [8]. R–U maps depict anonymized datasets according to their measured disclosure risk and data utility, facilitating the systematic comparison of alternative anonymization approaches. By anonymization approaches we refer to the different strategies by which datasets can be anonymized. These may involve method-specific variations, such as adjusting the level of noise in noise addition [1] or comparisons between methods – e.g., the effect of microaggregation [6] and post randomization [13] on disclosure risk. They can also include differences in synthetic data, either through employing distinct data generators and/or by producing multiple datasets with the same generator.
In practice, it is common to report one or more disclosure risk indicators and utility measures separately, often presented as summary statistics in tables or text, or visualized through simplified two-dimensional R–U maps [29, 43, 16, 23, 22]. Although such representations, especially R-U maps, aid in interpretation, displaying several risk and utility quickly becomes cumbersome. One remedy is to use faceting and arrange multiple R–U maps as small multiples: by enforcing comparisons of change, of differences between objects, small multiples are often the best solution for a wide range of presentation problems [47, p. 67-68]. However, as the number of measures grows and thereby the number of panels needed for their display, this approach becomes more fragmented and inefficient [17]. For example, with five distinct risk measures and five utility measures, 25 different R–U plots would be required to examine all pairwise relationships, complicating overall interpretation.
@@ 2.1 Pareto-Optimal/Efficient Trade-offs
The multivariate evaluation of disclosure risk and data utility naturally constitutes a multi-objective optimization problem, since improvements in one dimension often come at the expense of another. In such settings, the objectives are said to be at least partly conflicting [28, p. 5], and a solution (or observation) is called non-dominated, or Pareto-optimal/Pareto-efficient, if none of the objectives can be improved without degrading at least one of the others. Intuitively, this means that a Pareto-optimal solution represents an efficient trade-off: it cannot be improved in one aspect (e.g., utility) without performing worse in another (e.g., risk). Any method that is not Pareto-optimal is strictly inferior to at least one alternative [24, p. 547], a concept originating with Vilfredo Pareto’s Cours d’économie politique (1896–1897).
@@
with at least one strict inequality [28, p. 11] [9, Def. 5, p. 588]. An anonymization approach $i$ is Pareto-optimal if there exists no other anonymization approach $j$ that dominates it. This corresponds to strong Pareto optimality, since at least one inequality must be strict; the weaker notion, which also allows ties, is not considered here [28, p. 19].
@@
Additional decision criteria can also be applied. One option is to impose a risk-tolerance threshold and select the Pareto-optimal solution with the highest utility subject to this bound. Another intuitive “bang-for-buck” heuristic, applicable when the frontier is smooth and concave toward the origin, is to choose the knee (or elbow) point – where marginal increases in risk begin to yield only diminishing gains in utility, corresponding to the location of greatest curvature along the curve [45, p. 275].
@@ 2.2.3 Composite Risk/Utility scatterplots
Composite scatterplots summarizes the performance of different anonymization approaches by reducing multiple risk and utility measures into two composite scores [23, see, e.g.,]. The horizontal axis shows composite utility e.g. an average of all utility measures, while the vertical axis shows composite risk, similarly an average of all risk measures. Each point in the scatterplot corresponds to an single anonymization approach. The goal of the plot is to highlight trade-offs between disclosure risk and data utility, making it possible to identify anonymization strategies that strike a good balance. The quadrant structure is the same as in a classical R-U map and provides an intuitive interpretation (see Figure 3):
@@
While composite scores are commonly used to summarize performance across multiple risk or utility metrics, their interpretability hinges on the degree to which the constituent measures reflect a coherent underlying construct. To assess this, internal consistency metrics such as Cronbach’s $\alpha$ or McDonald’s $\omega$ can be used as heuristic diagnostics [3, 26, 56, 14]. A value of $\alpha\geq 0.70$ is often cited as indicative of acceptable consistency among items within a block [39, 31, p. 230]. A high value then means that the composite score is a reliable summary of the underlying set of risk or utility measures, reflecting shared variation rather than averaging over unrelated or inconsistent metrics. Note that the mentioned threshold is a rule-of-thumb and context-dependent, and $\alpha$ assumes tau-equivalence – i.e., equal contributions of all indicators to the latent construct – which may not hold in practice. McDonald’s $\omega$ is more general, as it allows heterogeneous loadings across items. In the context of multivariate risk–utility evaluation, internal consistency estimates serve only as rough checks for whether averaging across metrics (to form composite scores) is justifiable.
@@ 2.2.5 Radial Profile Plots (Radar / Origami)
Radial profile plots place each metric on a separate radial axis and connect the values for each anonymization approach or dataset to form a polygonal profile similar to PCP but on a circular basis (cf. Figure 6). Although originally proposed for variables with a circular course [25, p. 78], the visualization is now widely used for general multi-criteria comparison.
@@ 2.2.6 Multivariate PCA-based R-U Maps
    Biplots, introduced by [11], provide a graphical framework to represent both observations and variables in the same plot. By combining the PCA projection with variable loadings, biplots make it possible to explore patterns, relationships, and group structures in complex datasets. In the context of data anonymization, biplots provide an effective visualization for exploring the trade-off between disclosure risk and data utility while also revealing correlations among multiple evaluation metrics within a single, interpretable representation. Although dimensional reduction methods have previously been used to evaluate the utility of anonymized datasets [32, e.g.,], to the best of our knowledge, biplots have not yet been applied to assess multiple risk and utility measures simultaneously.
@@
    When outliers are a concern, a robust PCA variant (e.g., ROBPCA; [19]) is preferable. We further propose to use the score–distance / orthogonal-distance (SD-OD) diagnostic plot [19] to screen anonymization approaches for outliers. Following [19, p. 66], the robust score distance and orthogonal distance are
@@
    Region (SD, OD) Meaning Typical follow-up SD low, OD low Regular / typical; near the center of the data cloud and well represented by the first $k$ PCs. No concern; representative SDG. SD high, OD low Good Leverage outlier; far within the PC subspace (large scores on one or more PCs) yet small reconstruction error. Check influence; may be a valid extreme SDG. SD low, OD high Orthogonal outlier; not far in the PC plane but poorly reconstructed by the first $k$ PCs – structure outside the captured subspace. Inspect variables not explained by PCs; consider increasing $k$ or data issues. SD high, OD high Bad leverage; both far in the PC plane and poorly represented – extreme and structurally unusual. Strong candidate for anomaly; scrutinize or exclude.
    For a quick look, the same diagnostic plot can be made with classical PCA; for outlier detection and stable cutoffs we prefer robust PCA, comparing SD to $\sqrt{\chi^{2}_{k,.975}}$ and the OD to $\Phi^{-1}(.975)$ [19, p. 66]. Points beyond either cutoff are flagged as outliers cf. Table 1.
@@
    In the blockwise PCA approach, we conduct two separate principal component analyses: one on the set of disclosure risk measures and one on the set of utility measures [4, for utility measures only, see]. We then plot the first principal component of the utility block on the $x$-axis and the PC1 of the risk block on the $y$-axis. Each point in the resulting two-dimensional plot represents an anonymization approach, while the two axes serve as composite indices summarizing the respective blocks. Unlike the simple means used for the composites in the composite R-U map in Section 2.2.3, this PCA approach takes into account the correlation structure within each block and assigns weights to measures based on their contribution to overall variance. This allows for a more informed aggregation, where strongly varying and interrelated measures are emphasized, rather than treating all measures as equally important. However, this summarization comes at a cost: the axes are unitless and data-driven, and their interpretation depends on the specific loadings of the first principal components. To aid interpretation, the squared normalized loadings of PC1 can be visualized as stacked bar charts aligned to each axis (Utility PC1 and Risk PC1), showing which measures contribute most to the composite scores (cf. Figure 9). Optionally, Pareto-optimal anonymization approaches can be highlighted, and, when comparing multiple datasets for each anonymization approach, group structures can be visualized using ellipses to indicate clustering or method families.
@@ 3.1 Data and measures
Type Measure Abbreviation Description Risk Replicated Uniques [35] RepU Percentage of replicated sample uniques in synthetic dataset. Risk Disclosure by Successful Correct Outcome [35] DiSCO Percentage of disclosive in synthetic and correct in the original. Risk Disclosure by Correct Attribution Probability (Direct) [41] DCAP Probability that a synthetic record can be correctly attributed to an original record (direct case). Risk Targeted Correct Attribution Probability [40] TCAP Probability of correct attribution when targeting specific records. Risk Risk of Attribute Prediction-Induced Disclosure (Currently under development by the authors; a preprint will be available within the next month.) RAPID Expected share of re-identifiable records. Utility Confidence Interval Proximity Proximity Deviation of confidence intervals for sensitive variables. Utility Propensity Mean Squared Error [37] pMSE Predictive score from distinguishing real vs. synthetic data. Utility Wasserstein Distance [49] Wasserstein Wasserstein distance between numeric distributions. Utility Hellinger Distance [15] Hellinger Hellinger distance between categorical distributions. Utility Energy Distance [36] Energy Multivariate energy distance on numeric variables.
@@ 3.2 Visualization Approaches
Figure 8: Robust PCA diagnostics (“outlier map”). The x-axis shows the Score Distance (SD) in the retained PC space (spanned by $k=2$ components), and the y-axis shows the Orthogonal Distance (OD) to that space. Each point is an SDG; color indicates the robust outlier flag. Large SD indicates leverage within the PC space; large OD indicates poor reconstruction (the observation lies far outside the subspace). Cutoffs follow [19, p. 66 ]: SD is compared to $\sqrt{\chi^{2}_{k,0.975}}$ and OD to $\Phi^{-1}(0.975)$.
@@ Appendix A Synthetic Data Generators
Name Availability Method(s) Software Year Last upd. synthpop [^1] https://cran.r-project.org/web/packages/synthpop/ (accessed 2024-06-26) open source CART R package 2014 2025 Synthetic Data Vault [^2] https://pypi.org/project/sdv/0.3.1/ (accessed 2025-22-09) open source GaussianCopula; CTGAN; VAE Python package 2018 2025 simPop [^3] https://cran.r-project.org/web/packages/simPop/ (accessed 2025-22-09) open source Multinomial log-linear models; random draws; random forest (alt.) R package 2010 2024 Mostly AI [^4] https://mostly.ai/ (accessed 2025-22-09) proprietary Transformers; GANs; VAEs; autoregressive networks Mostly AI (AT) 2017 2024 Gretel [^5] https://gretel.ai/ (accessed 2025-22-09) both Synthetic ACTGAN Gretel Labs (US) 2020 2024 arf [^6] https://cran.r-project.org/web/packages/arf/ (accessed 2025-22-09) open source Advanced random forests R package 2022 2025 synthcity [^7] https://pypi.org/project/synthcity/ (accessed 2025-22-09) open source Advanced random forests R package 2022 2025

### v2

@@ Abstract
Keywords: statistical disclosure control, risk-utility, Pareto optimality, multivariate statistics, visualization
@@ 1 Introduction
In microdata anonymization, modifications to the original dataset, such as suppression, generalization, perturbation, or synthetic data generation, are essential to reduce disclosure risk [21, 47]. However, these modifications inevitably result in a loss of information, potentially limiting the analytical value of the data. Effective anonymization therefore requires balancing the confidentiality gained from lowering disclosure risk with the preservation of data utility. A possibility for evaluating this trade-off visually is the risk–utility (R-U) confidentiality map [8]. R–U maps depict anonymized datasets according to their measured disclosure risk and data utility, facilitating the systematic comparison of alternative anonymization approaches. By anonymization approaches, we refer to the different strategies by which datasets can be anonymized. These may involve method-specific variations, such as adjusting the level of noise in noise addition [1] or comparisons between methods – e.g., the effect of microaggregation [6] and post randomization [13] on disclosure risk. They can also include differences in synthetic data, either through employing distinct data generators and/or by producing multiple datasets with the same generator.
In practice, it is common to report one or more disclosure risk indicators and utility measures separately, often presented as summary statistics in tables or text, or visualized through simplified two-dimensional R–U maps [31, 48, 17, 24, 23]. Although such representations, especially R-U maps, aid in interpretation, displaying several risk and utility quickly becomes cumbersome. One remedy is to use faceting and arrange multiple R–U maps as small multiples: by enforcing comparisons of change, of differences between objects, small multiples are often the best solution for a wide range of presentation problems [52, p. 67-68]. However, as the number of measures grows and thereby the number of panels needed for their display, this approach becomes more fragmented and inefficient [18]. For example, with five distinct risk measures and five utility measures, 25 different R–U plots would be required to examine all pairwise relationships, complicating overall interpretation.
@@ 2.1 Pareto-Optimal/Efficient Trade-offs
The multivariate evaluation of disclosure risk and data utility naturally constitutes a multi-objective optimization problem, since improvements in one dimension often come at the expense of another. In such settings, the objectives are said to be at least partly conflicting [29, p. 5], and a solution (or observation) is called non-dominated, or Pareto-optimal/Pareto-efficient, if none of the objectives can be improved without degrading at least one of the others. Intuitively, this means that a Pareto-optimal solution represents an efficient trade-off: it cannot be improved in one aspect (e.g., utility) without performing worse in another (e.g., risk). Any method that is not Pareto-optimal is strictly inferior to at least one alternative [25, p. 547], a concept originating with Vilfredo Pareto’s Cours d’économie politique (1896–1897).
@@
with at least one strict inequality [29, p. 11] [9, Def. 5, p. 588]. An anonymization approach $i$ is Pareto-optimal if there exists no other anonymization approach $j$ that dominates it. This corresponds to strong Pareto optimality, since at least one inequality must be strict; the weaker notion, which also allows ties, is not considered here [29, p. 19].
@@
Additional decision criteria can also be applied. One option is to impose a risk-tolerance threshold and select the Pareto-optimal solution with the highest utility subject to this bound. Another intuitive “bang-for-buck” heuristic, applicable when the frontier is smooth and concave toward the origin, is to choose the knee (or elbow) point – where marginal increases in risk begin to yield only diminishing gains in utility, corresponding to the location of greatest curvature along the curve [50, p. 275].
@@ 2.2.3 Composite Risk/Utility scatterplots
Composite scatterplots summarizes the performance of different anonymization approaches by reducing multiple risk and utility measures into two composite scores [[, see, e.g.,]]Little_2022_Comparing. The horizontal axis shows composite utility e.g. an average of all utility measures, while the vertical axis shows composite risk, similarly an average of all risk measures. Each point in the scatterplot corresponds to an single anonymization approach. The goal of the plot is to highlight trade-offs between disclosure risk and data utility, making it possible to identify anonymization strategies that strike a good balance. The quadrant structure is the same as in a classical R-U map and provides an intuitive interpretation (see Figure 3):
@@
While composite scores are commonly used to summarize performance across multiple risk or utility metrics, their interpretability hinges on the degree to which the constituent measures reflect a coherent underlying construct. To assess this, internal consistency metrics such as Cronbach’s $\alpha$ or McDonald’s $\omega$ can be used as heuristic diagnostics [3, 27, 62, 15]. A value of $\alpha\geq 0.70$ is often cited as indicative of acceptable consistency among items within a block [44, 34, p. 230]. A high value then means that the composite score is a reliable summary of the underlying set of risk or utility measures, reflecting shared variation rather than averaging over unrelated or inconsistent metrics. Note that the mentioned threshold is a rule-of-thumb and context-dependent, and $\alpha$ assumes tau-equivalence – i.e., equal contributions of all indicators to the latent construct – which may not hold in practice. McDonald’s $\omega$ is more general, as it allows heterogeneous loadings across items. In the context of multivariate risk–utility evaluation, internal consistency estimates serve only as rough checks for whether averaging across metrics (to form composite scores) is justifiable.
@@ 2.2.5 Radial Profile Plots (Radar / Origami)
Radial profile plots place each metric on a separate radial axis and connect the values for each anonymization approach or dataset to form a polygonal profile similar to PCP but on a circular basis (cf. Figure 6). Although originally proposed for variables with a circular course [26, p. 78], the visualization is now widely used for general multi-criteria comparison.
@@ 2.2.6 Multivariate PCA-based R-U Maps
    Biplots, introduced by [11], provide a graphical framework to represent both observations and variables in the same plot. By combining the PCA projection with variable loadings, biplots make it possible to explore patterns, relationships, and group structures in complex datasets. In the context of data anonymization, biplots provide an effective visualization for exploring the trade-off between disclosure risk and data utility while also revealing correlations among multiple evaluation metrics within a single, interpretable representation. Although dimensional reduction methods have previously been used to evaluate the utility of anonymized datasets [[, e.g.,]]2025_Bachot, to the best of our knowledge, biplots have not yet been applied to assess multiple risk and utility measures simultaneously.
@@
    When outliers are a concern, a robust PCA variant (e.g., ROBPCA; [20]) is preferable. We further propose to use the score–distance / orthogonal-distance (SD-OD) diagnostic plot [20] to screen anonymization approaches for outliers. Following [20, p. 66], the robust score distance and orthogonal distance are
@@
    Region (SD, OD) Meaning Typical follow-up
    SD low, OD low Regular / typical; near the center of the data cloud and well represented by the first $k$ PCs. No concern; representative SDG.
    SD high, OD low Good Leverage outlier; far within the PC subspace (large scores on one or more PCs) yet small reconstruction error. Check influence; may be a valid extreme SDG.
    SD low, OD high Orthogonal outlier; not far in the PC plane but poorly reconstructed by the first $k$ PCs – structure outside the captured subspace. Inspect variables not explained by PCs; consider increasing $k$ or data issues.
    SD high, OD high Bad leverage; both far in the PC plane and poorly represented – extreme and structurally unusual. Strong candidate for anomaly; scrutinize or exclude.
    For a quick look, the same diagnostic plot can be made with classical PCA; for outlier detection and stable cutoffs we prefer robust PCA, comparing SD to $\sqrt{\chi^{2}_{k,.975}}$ and the OD to $\Phi^{-1}(.975)$ [20, p. 66]. Points beyond either cutoff are flagged as outliers cf. Table 1.
@@
    In the blockwise PCA approach, we conduct two separate principal component analyses: one on the set of disclosure risk measures and one on the set of utility measures [[, for utility measures only, see]]2022_Dankar_PCA. We then plot the first principal component of the utility block on the $x$-axis and the PC1 of the risk block on the $y$-axis. Each point in the resulting two-dimensional plot represents an anonymization approach, while the two axes serve as composite indices summarizing the respective blocks. Unlike the simple means used for the composites in the composite R-U map in Section 2.2.3, this PCA approach takes into account the correlation structure within each block and assigns weights to measures based on their contribution to overall variance. This allows for a more informed aggregation, where strongly varying and interrelated measures are emphasized, rather than treating all measures as equally important. However, this summarization comes at a cost: the axes are unitless and data-driven, and their interpretation depends on the specific loadings of the first principal components. To aid interpretation, the squared normalized loadings of PC1 can be visualized as stacked bar charts aligned to each axis (Utility PC1 and Risk PC1), showing which measures contribute most to the composite scores (cf. Figure 9). Optionally, Pareto-optimal anonymization approaches can be highlighted, and, when comparing multiple datasets for each anonymization approach, group structures can be visualized using ellipses to indicate clustering or method families.
@@ 3.1 Data and measures
Type Measure & Abbreviation Author Description Risk Replicated Uniques (RepU ) [40] Percentage of replicated sample uniques in synthetic dataset. Risk Disclosure by Successful Correct Outcome (DiSCO) [40] Percentage of disclosive in synthetic and correct in the original. Risk Disclosure by Correct Attribution Probability (Direct) (DCAP) [46] Probability that a synthetic record can be correctly attributed to an original record (direct case). Risk Targeted Correct Attribution Probability (TCAP ) [45] Probability of correct attribution when targeting specific records. Risk Risk of Attribute Prediction-Induced Disclosure (RAPID) (Currently under development by the authors; a preprint will be available within the next month.) Expected share of re-identifiable records. Utility Confidence Interval Proximity (Proximity) Deviation of confidence intervals for sensitive variables. Utility Propensity Mean Squared Error (pMSE ) [42] Predictive score from distinguishing real vs. synthetic data. Utility Wasserstein Distance (Wasserstein) [54] Wasserstein distance between numeric distributions. Utility Hellinger Distance (Hellinger) [16] Hellinger distance between categorical distributions. Utility Energy Distance (Energy) [41] Multivariate energy distance on numeric variables.
@@ 3.2 Visualization Approaches
Figure 8: Robust PCA diagnostics (“outlier map”). The x-axis shows the Score Distance (SD) in the retained PC space (spanned by $k=2$ components), and the y-axis shows the Orthogonal Distance (OD) to that space. Each point is an SDG; color indicates the robust outlier flag. Large SD indicates leverage within the PC space; large OD indicates poor reconstruction (the observation lies far outside the subspace). Cutoffs follow [20, p. 66]: SD is compared to $\sqrt{\chi^{2}_{k,0.975}}$ and OD to $\Phi^{-1}(0.975)$.
@@ Appendix A Synthetic Data Generators
Name Method(s) Software Authors Version synthpop CART R package [33] 1.9.1 Synthetic Data Vault GaussianCopula; VAE Python package [35] 1.18.0 simPop Multinomial log-linear models; random draws; random forest (alt.) R package [49] 2.1.3 Mostly AI Transformers; GANs; VAEs; autoregressive networks Mostly AI (AT) [30] 4.2.3 Gretel Synthetic ACTGAN Gretel Labs (US) [14] 0.22.16 arf Advanced random forests R package [57] 0.2.0 synthcity Tabular GAN Python package [38] 0.2.11